# GRAND: Graph Release with Assured Node Differential Privacy

### v2

@@ B.3 Latent distribution consistency: multidimensional case
    $|\hat{F}_{2|1,m}(\hat{Z}_{i2}|\hat{Z}_{i1})-\hat{F}_{2|1,m}(Z_{i2}|\hat{Z}_{i1})|\leq L_{\hat{F}}|\hat{Z}_{i2}-Z_{i2}|=o_{P}(1).$
  • Term 2: This term relies on $\hat{F}_{2|1,m}(x|u)$ being uniformly equicontinuous with respect to its conditioning argument $u$. So, $|\hat{F}_{2|1,m}(Z_{i2}|\hat{Z}_{i1})-\hat{F}_{2|1,m}(Z_{i2}|Z_{i1})|\xrightarrow{P}0$ because $\|\hat{Z}_{i1}-Z_{i1}\|\leq\delta_{m}\to 0$.
These indicate $|q_{A}-q_{B}|\xrightarrow{P}0$.
As discussed, this proves the convergence of $\tilde{Z}_{i12}$ (conditioning on $\tilde{Z}_{i1}$).
@@ Appendix C Proof of Theorem 3 (Emprical CDF Convergence)
Again, we first work on the univariate version, $d=1$. And we start with the intermediate random vector $\bar{Z}_{i}$ again. Suppose $d=1$. Define the empirical CDF of $\bar{Z}_{i}$’s as
$\bar{F}_{n}(x)=\frac{1}{n}\sum_{1\leq i\leq n}\mathbf{1}_{\{\bar{Z}_{i}\leq x\}}$
and
$\tilde{F}_{n}(x)=\frac{1}{n}\sum_{1\leq i\leq n}\mathbf{1}_{\{\tilde{Z}_{i}\leq x\}}$
Lemma C.1.
@@
$\sup_{x}|\bar{F}_{n}(x)-F(x)|\xrightarrow{P}0$
as $m,n\to\infty$.
@@
$|\bar{F}_{n}(x)-F_{n}(x)|$ $=|\frac{1}{n}\sum_{1\leq i\leq n}\mathbf{1}_{\{\bar{Z}_{i}\leq x\}}-\frac{1}{n}\sum_{1\leq i\leq n}\mathbf{1}_{\{T_{i}\leq x\}}|$
$\leq\frac{1}{n}\sum_{i}|\mathbf{1}_{\{\bar{Z}_{i}\leq x\}}-\mathbf{1}_{\{T_{i}\leq x\}}|.$
@@
$\sup_{x}|\bar{F}_{n}(x)-F_{n}(x)|$ $\leq\sup_{x}\frac{1}{n}\sum_{i}|\mathbf{1}_{\{\bar{Z}_{i}\leq x\}}-\mathbf{1}_{\{T_{i}\leq x\}}|$
$\leq\sup_{x}\big{\{}\frac{1}{n}\sum_{i:|\bar{Z}_{i}-T_{i}|<\varepsilon}|\mathbf{1}_{\{\bar{Z}_{i}\leq x\}}-\mathbf{1}_{\{T_{i}\leq x\}}|+\frac{1}{n}\sum_{i:|\bar{Z}_{i}-T_{i}|\geq\varepsilon}|\mathbf{1}_{\{\bar{Z}_{i}\leq x\}}-\mathbf{1}_{\{T_{i}\leq x\}}|\big{\}}$
$\leq\sup_{x}\big{\{}\frac{1}{n}\sum_{i:|\bar{Z}_{i}-T_{i}|<\varepsilon}\mathbf{1}_{\{T_{i}\in B(x,\varepsilon)\}}+\frac{1}{n}\#\{i:|\bar{Z}_{i}-T_{i}|\geq\varepsilon\}\big{\}}$
$\leq\sup_{x}\big{\{}\frac{1}{n}\sum_{i}\mathbf{1}_{\{T_{i}\in B(x,\varepsilon)\}}+\frac{1}{n}\#\{i:|\bar{Z}_{i}-T_{i}|\geq\varepsilon\}\big{\}}$
$\leq O_{P,n}(\varepsilon)+\frac{1}{n}\sum_{i}\mathbf{1}_{|\{\bar{Z}_{i}-T_{i}|\geq\varepsilon\}}.$ (20)
where $O_{P,n}$ denotes that quantity order with high probability in $n$.
@@
$A_{i}=\hat{F}_{m}^{-1}\left(G\left(\hat{F}_{m}(Z_{i})+e_{i}\right)\right)-\hat{F}_{m}^{-1}\left(G\left(F(Z_{i})+e_{i}\right)\right),$
$B_{i}=\hat{F}_{m}^{-1}\left(G\left(F(Z_{i})+e_{i}\right)\right)-F^{-1}\left(G\left(F(Z_{i})+e_{i}\right)\right).$
@@
Recall that in that context, we take an interval $[a,b]$, and here we specify $a,b$ to be the $\varepsilon/16$ and $1-\varepsilon/16$ quantiles of $U(0,1)$. We have defined
$\mathscr{E}_{m}=\left\{\sup_{1\leq i\leq n}|\hat{F}_{m}(Z_{i})-F(Z_{i})|<\delta\right\}\cap\left\{\sup_{\begin{subarray}{c}u,v\in[a,b]\\ |u-v|<\delta^{\prime}\end{subarray}}|\hat{F}_{m}^{-1}(u)-\hat{F}_{m}^{-1}(v)|<\varepsilon/2\right\}$
We have seen that $P(\mathscr{E}_{m})\to 1$ and on $\mathscr{E}_{m}$, we have $\sup_{i}|A_{i}|<\varepsilon/2$.
Unfortunately, the other event in the proof, $\mathscr{E}_{m}^{\prime}$ depends on $i$, so we could not achieve the uniform control on $B_{i}$’s. Instead, let us define $\mathscr{E}_{m}^{\prime}=\left\{\sup_{u\in[a,b]}|\hat{F}_{m}^{-1}(u)-F^{-1}(u)|<\varepsilon/2\right\}$. Again, $P(\mathscr{E}_{m}^{\prime})\to 1$ as shown in Lemma B.3.
@@
$\frac{1}{n}\sum_{i}\mathbf{1}_{|\{\bar{Z}_{i}-T_{i}|\geq\varepsilon\}}\leq\frac{1}{n}\sum_{i}\mathbf{1}_{|\{U_{i}\notin[a,b]\}}.$
By Hoefding’s inequality, we can further define an event $\mathscr{E}_{m}^{{}^{\prime\prime}}$ such that
$\frac{1}{n}\sum_{i}\mathbf{1}_{|\{U_{i}\notin[a,b]\}}\leq\frac{\varepsilon}{2}$
@@
$\sup_{x}|\tilde{F}_{n}(x)-F(x)|\xrightarrow{P}0$
as $m,n\to\infty$.
@@
$\sup_{x}|\tilde{F}_{n}(x)-\bar{F}_{n}(x)|\xrightarrow{P}0.$
@@
$\sup_{x}|\bar{F}_{n}(x)-\tilde{F}_{n}(x)|$ $\leq\sup_{x}\frac{1}{n}\sum_{i}|\mathbf{1}_{\{\bar{Z}_{i}\leq x\}}-\mathbf{1}_{\{\tilde{Z}_{i}\leq x\}}|$
$\leq\sup_{x}\big{\{}\frac{1}{n}\sum_{i:|\bar{Z}_{i}-\tilde{Z}_{i}|<\varepsilon}|\mathbf{1}_{\{\bar{Z}_{i}\leq x\}}-\mathbf{1}_{\{\tilde{Z}_{i}\leq x\}}|+\frac{1}{n}\sum_{i:|\bar{Z}_{i}-\tilde{Z}_{i}|\geq\varepsilon}|\mathbf{1}_{\{\bar{Z}_{i}\leq x\}}-\mathbf{1}_{\{\tilde{Z}_{i}\leq x\}}|\big{\}}$
$\leq\sup_{x}\big{\{}\frac{1}{n}\sum_{i:|\bar{Z}_{i}-\tilde{Z}_{i}|<\varepsilon}\mathbf{1}_{\{\bar{Z}_{i}\in B(x,\varepsilon)\}}+\frac{1}{n}\sum_{i}\mathbf{1}_{\{|\bar{Z}_{i}-\tilde{Z}_{i}|\geq\varepsilon\}}\big{\}}$
$\leq\sup_{x}\big{\{}\frac{1}{n}\sum_{i}\mathbf{1}_{\{\bar{Z}_{i}\in B(x,\varepsilon)\}}+\frac{1}{n}\sum_{i}\mathbf{1}_{\{|\bar{Z}_{i}-\tilde{Z}_{i}|\geq\varepsilon\}}\big{\}}.$ (21)
Note that $\tilde{Z}_{i},i\in[n]$ are independent, because $\hat{F}_{m}$ is based only on $Z_{i},i>n$. For the first term $\frac{1}{n}\sum_{i}\mathbf{1}_{\{\bar{Z}_{i}\in B(x,\varepsilon)\}}$, note that $P(Z_{i}\in B(x,\varepsilon))=O(\varepsilon)$ in the current assumption and because of the uniform convergence of $\bar{F}_{n}$ and the independence between $\tilde{Z}_{i}$’s, we have
$\sup_{x}\frac{1}{n}\sum_{i}\mathbf{1}_{\{\bar{Z}_{i}\in B(x,\varepsilon)\}}=O_{P,n}(\varepsilon).$
@@
$u_{i}=G\left(\hat{F}_{m}(\hat{Z}_{i})+e_{i}\right),\quad v_{i}=G\left(\hat{F}_{m}(Z_{i})+e_{i}\right).$
For the current $\epsilon$, pick $a,b$ such that $P(U\sim[a,b])\geq 1-\varepsilon/4$. We have
$\big{\{}i:|\bar{Z}_{i}-\tilde{Z}_{i}|\geq\varepsilon\big{\}}$
$\subset$ $\big{\{}i:|\bar{Z}_{i}-\tilde{Z}_{i}|\geq\varepsilon,v_{i},u_{i}\in[a,b]\big{\}}\cup\big{\{}i:v_{i}\notin[a,b]\big{\}}\cup\big{\{}i:u_{i}\notin[a,b]\big{\}}$
$=$ $\big{\{}i:|\hat{F}_{m}^{-1}(u_{i})-\hat{F}_{m}^{-1}(v_{i})|\geq\varepsilon,v_{i},u_{i}\in[a,b]\big{\}}\cup\big{\{}i:v_{i}\notin[a,b]\big{\}}\cup\big{\{}i:u_{i}\notin[a,b]\big{\}}$
@@
$\frac{1}{n}\sum_{i}\mathbf{1}_{\{|\bar{Z}_{i}-T_{i}|\geq\varepsilon\}}\leq\frac{1}{n}\sum_{i}\mathbf{1}_{\{|\hat{F}_{m}^{-1}(u_{i})-\hat{F}_{m}^{-1}(v_{i})|\geq\varepsilon,v_{i},u_{i}\in[a,b]\}}+\frac{1}{n}\sum_{i}\mathbf{1}_{v_{i}\notin[a,b]}+\frac{1}{n}\sum_{i}\mathbf{1}_{u_{i}\notin[a,b]}.$
As the proof of Lemma B.4, Theorem B.1 indicates that there exists $\delta>0$ and $M>0$, such that for any $m>M$,
$\sup_{u,v\in[a,b],|u-v|<\delta}|\hat{F}^{-1}_{m}(u)-\hat{F}^{-1}_{m}(v)|<\epsilon$
with probability larger than $1-\epsilon/4.$ Define the event $\mathscr{E}^{{}^{\prime\prime\prime}}$ to be the intersect of this event and the event $\sup_{i}|G(u_{i})-G(v_{i})|<\delta$. Under the current assumptions of $\hat{Z}_{i}$’s concentration and Lemma B.1, we can see that $P(\mathscr{E}^{{}^{\prime\prime\prime}})\to 1.$
@@
$u_{i}-G(F(Z_{i}))\xrightarrow{P}0,v_{i}-G(F(Z_{i}))\xrightarrow{P}0.$
Additionally, since $G(F(Z_{i}))\sim U(0,1)$, we have
$\frac{1}{n}\sum_{i}\mathbf{1}_{v_{i}\notin[a,b]}=O_{P,m,n}(\epsilon),\quad\frac{1}{n}\sum_{i}\mathbf{1}_{u_{i}\notin[a,b]}=O_{P,m,n}(\epsilon).$
Combining these, we have $\frac{1}{n}\sum_{i}\mathbf{1}_{\{|\bar{Z}_{i}-T_{i}|\geq\varepsilon\}}=O(P,m,n)(\epsilon)$.
Because $\epsilon$ is arbitrary, we complete the proof.
@@
$\sup_{(x_{1},x_{2})\in{\bf R}^{2}}|\bar{F}_{n}(x_{1},x_{2})-F(x_{1},x_{2})|\xrightarrow{P}0$
@@
By construction, $T_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}F$ , so by the multivariate Glivenko-Cantelli theorem:
$\sup_{(x_{1},x_{2})\in{\bf R}^{2}}|F_{n}(x_{1},x_{2})-F(x_{1},x_{2})|\xrightarrow{P}0.$
@@
$\sup_{(x_{1},x_{2})\in{\bf R}^{2}}|\bar{F}_{n}(x_{1},x_{2})-F_{n}(x_{1},x_{2})|\xrightarrow{P}0.$
For any $(x_{1},x_{2})$:
$|\bar{F}_{n}(x_{1},x_{2})-F_{n}(x_{1},x_{2})|$ $=\Big{|}\frac{1}{n}\sum_{i=1}^{n}[\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1},\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1},T_{i2}\leq x_{2}\}}]\Big{|}$

### v3

@@ B.3 Latent distribution consistency: multidimensional case
    $\big|\hat{F}_{2\,|\,1,m}(\hat{Z}_{i2}\,|\,\hat{Z}_{i1})-\hat{F}_{2\,|\,1,m}(Z_{i2}\,|\,\hat{Z}_{i1})\big|\leq L_{\hat{F}_{2\,|\,1,m}}|\hat{Z}_{i2}-Z_{i2}|=o_{\mathbb{P}}(1).$
  • Term 2: This term relies on $\hat{F}_{2\,|\,1,m}(x\,|\,u)$ being uniformly equicontinuous with respect to its conditioning argument $u$. So, $\big|\hat{F}_{2\,|\,1,m}(Z_{i2}\,|\,\hat{Z}_{i1})-\hat{F}_{2\,|\,1,m}(Z_{i2}\,|\,Z_{i1})\big|\xrightarrow{\mathbb{P}}0$ because $|\hat{Z}_{i1}-Z_{i1}|\leq\delta_{m}\xrightarrow{\mathbb{P}}0$.
These indicate $|q_{A}-q_{B}|\xrightarrow{\mathbb{P}}0$.
As discussed, this proves the convergence of $\tilde{Z}_{i2}$ (conditioning on $\tilde{Z}_{i1}$).
@@ Appendix C Proof of Theorem 3 (Emprical CDF Convergence)
Again, we first work on the univariate version, $d=1$. And we start with the intermediate random vector $\bar{Z}_{i}$ again. Suppose $d=1$. Define the empirical CDFs of $\bar{Z}_{i}$’s and $\tilde{Z}_{i}$’s as
$\bar{F}_{n}(x)=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{\bar{Z}_{i}\leq x\}},\quad\tilde{F}_{n}(x)=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{\tilde{Z}_{i}\leq x\}}.$
Lemma C.1 (One-dimensional Intermediate Convergence).
@@
$\sup_{x\in\mathbb{R}}\big|\bar{F}_{n}(x)-F(x)\big|\xrightarrow{\mathbb{P}}0\quad\text{as }m,n\to\infty.$
@@
$\big|\bar{F}_{n}(x)-F_{n}(x)\big|$ $=\left|\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{\bar{Z}_{i}\leq x\}}-\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{T_{i}\leq x\}}\right|$
$\leq\frac{1}{n}\sum_{i=1}^{n}\big|\mathbf{1}_{\{\bar{Z}_{i}\leq x\}}-\mathbf{1}_{\{T_{i}\leq x\}}\big|.$
@@
$\sup_{x\in\mathbb{R}}\big|\bar{F}_{n}(x)-F_{n}(x)\big|$ $\leq\sup_{x\in\mathbb{R}}\frac{1}{n}\sum_{i=1}^{n}\big|\mathbf{1}_{\{\bar{Z}_{i}\leq x\}}-\mathbf{1}_{\{T_{i}\leq x\}}\big|$
$\leq\sup_{x\in\mathbb{R}}\left\{\frac{1}{n}\sum_{i:|\bar{Z}_{i}-T_{i}|<\varepsilon}|\mathbf{1}_{\{\bar{Z}_{i}\leq x\}}-\mathbf{1}_{\{T_{i}\leq x\}}|+\frac{1}{n}\sum_{i:|\bar{Z}_{i}-T_{i}|\geq\varepsilon}|\mathbf{1}_{\{\bar{Z}_{i}\leq x\}}-\mathbf{1}_{\{T_{i}\leq x\}}|\right\}$
$\leq\sup_{x\in\mathbb{R}}\left\{\frac{1}{n}\sum_{i:|\bar{Z}_{i}-T_{i}|<\varepsilon}\mathbf{1}_{\{T_{i}\in B(x,\varepsilon)\}}+\frac{1}{n}\#\{i:|\bar{Z}_{i}-T_{i}|\geq\varepsilon\}\right\}$
$\leq\sup_{x\in\mathbb{R}}\left\{\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{T_{i}\in B(x,\varepsilon)\}}+\frac{1}{n}\#\{i:|\bar{Z}_{i}-T_{i}|\geq\varepsilon\}\right\}$
$\leq\mathcal{O}_{\mathbb{P},n}(\varepsilon)+\frac{1}{n}\sum_{i}\mathbf{1}_{\{|\bar{Z}_{i}-T_{i}|\geq\varepsilon\}}.$ (20)
where $\mathcal{O}_{\mathbb{P},n}$ denotes that quantity order with high probability in $n$.
@@
$A_{i}=\hat{F}_{m}^{-1}(G(\hat{F}_{m}(Z_{i})+e_{i}))-\hat{F}_{m}^{-1}(G(F(Z_{i})+e_{i})),$
$B_{i}=\hat{F}_{m}^{-1}(G(F(Z_{i})+e_{i}))-F^{-1}(G(F(Z_{i})+e_{i})).$
@@
Recall that in that context, we take an interval $[a,b]$, and here we specify $a,b$ to be the $\varepsilon/16$ and $1-\varepsilon/16$ quantiles of $\text{Uniform}(0,1)$. We have defined
$\mathscr{E}_{m}=\left\{\sup_{1\leq i\leq n}\big|\hat{F}_{m}(Z_{i})-F(Z_{i})\big|<\delta\right\}\cap\left\{\sup_{\begin{subarray}{c}u,v\in[a,b]\\ |u-v|<\delta^{\prime}\end{subarray}}\big|\hat{F}_{m}^{-1}(u)-\hat{F}_{m}^{-1}(v)\big|<\frac{\varepsilon}{2}\right\}.$
We have seen that $\mathbb{P}(\mathscr{E}_{m})\xrightarrow{m\to\infty}1$ and on $\mathscr{E}_{m}$, we have $\sup_{i\in[n]}|A_{i}|<\varepsilon/2$.
Unfortunately, the other event in the proof, $\mathscr{E}_{m}^{\prime}$ depends on $i$, so we could not achieve the uniform control on $B_{i}$’s. Instead, let us define $\mathscr{E}_{m}^{\prime}=\left\{\sup_{u\in[a,b]}\big|\hat{F}_{m}^{-1}(u)-F^{-1}(u)\big|<\varepsilon/2\right\}$. Again, $\mathbb{P}(\mathscr{E}_{m}^{\prime})\xrightarrow{m\to\infty}1$ as shown in Lemma B.3.
@@
$\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{|\bar{Z}_{i}-T_{i}|\geq\varepsilon\}}\leq\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{U_{i}\notin[a,b]\}}.$
By Hoeffding’s inequality, we can further define an event $\mathscr{E}_{m}^{{}^{\prime\prime}}$ such that
$\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{U_{i}\notin[a,b]\}}\leq\frac{\varepsilon}{2}$
@@
$\sup_{x\in\mathbb{R}}\big|\tilde{F}_{n}(x)-F(x)\big|\xrightarrow{\mathbb{P}}0\quad\text{as }m,n\to\infty.$
@@
$\sup_{x\in\mathbb{R}}\big|\tilde{F}_{n}(x)-\bar{F}_{n}(x)\big|\xrightarrow{\mathbb{P}}0.$
@@
$\sup_{x\in\mathbb{R}}\big|\bar{F}_{n}(x)-\tilde{F}_{n}(x)\big|$ $\leq\sup_{x\in\mathbb{R}}\frac{1}{n}\sum_{i=1}^{n}\big|\mathbf{1}_{\{\bar{Z}_{i}\leq x\}}-\mathbf{1}_{\{\tilde{Z}_{i}\leq x\}}\big|$
$\leq\sup_{x\in\mathbb{R}}\left\{\frac{1}{n}\sum_{i:|\bar{Z}_{i}-\tilde{Z}_{i}|<\varepsilon}\big|\mathbf{1}_{\{\bar{Z}_{i}\leq x\}}-\mathbf{1}_{\{\tilde{Z}_{i}\leq x\}}\big|+\frac{1}{n}\sum_{i:|\bar{Z}_{i}-\tilde{Z}_{i}|\geq\varepsilon}\big|\mathbf{1}_{\{\bar{Z}_{i}\leq x\}}-\mathbf{1}_{\{\tilde{Z}_{i}\leq x\}}\big|\right\}$
$\leq\sup_{x\in\mathbb{R}}\left\{\frac{1}{n}\sum_{i:|\bar{Z}_{i}-\tilde{Z}_{i}|<\varepsilon}\mathbf{1}_{\{\bar{Z}_{i}\in B(x,\varepsilon)\}}+\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{|\bar{Z}_{i}-\tilde{Z}_{i}|\geq\varepsilon\}}\right\}$
$\leq\sup_{x\in\mathbb{R}}\left\{\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{\bar{Z}_{i}\in B(x,\varepsilon)\}}+\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{|\bar{Z}_{i}-\tilde{Z}_{i}|\geq\varepsilon\}}\right\}.$ (21)
Note that $\tilde{Z}_{i}$, $i\in[n]$ are independent, because $\hat{F}_{m}$ is based only on $Z_{i}$, $i>n$. For the first term $\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{\bar{Z}_{i}\in B(x,\varepsilon)\}}$, note that $\mathbb{P}(Z_{i}\in B(x,\varepsilon))=\mathcal{O}(\varepsilon)$ in the current assumption and because of the uniform convergence of $\bar{F}_{n}$ and the independence between $\tilde{Z}_{i}$’s, we have
$\sup_{x\in\mathbb{R}}\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{\bar{Z}_{i}\in B(x,\varepsilon)\}}=\mathcal{O}_{\mathbb{P},n}(\varepsilon).$
@@
$u_{i}=G(\hat{F}_{m}(\hat{Z}_{i})+e_{i}),\quad v_{i}=G(\hat{F}_{m}(Z_{i})+e_{i}).$
For the current $\varepsilon$, pick $a,b$ such that $\mathbb{P}(U\in[a,b])\geq 1-\varepsilon/4$. We have
$\left\{i:|\bar{Z}_{i}-\tilde{Z}_{i}|\geq\varepsilon\right\}$
$\subset$ $\left\{i:|\bar{Z}_{i}-\tilde{Z}_{i}|\geq\varepsilon,v_{i},u_{i}\in[a,b]\right\}\cup\left\{i:v_{i}\notin[a,b]\right\}\cup\left\{i:u_{i}\notin[a,b]\right\}$
$=$ $\left\{i:\big|\hat{F}_{m}^{-1}(u_{i})-\hat{F}_{m}^{-1}(v_{i})\big|\geq\varepsilon,v_{i},u_{i}\in[a,b]\right\}\cup\left\{i:v_{i}\notin[a,b]\right\}\cup\left\{i:u_{i}\notin[a,b]\right\}.$
@@
$\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{|\bar{Z}_{i}-T_{i}|\geq\varepsilon\}}\leq\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{|\hat{F}_{m}^{-1}(u_{i})-\hat{F}_{m}^{-1}(v_{i})|\geq\varepsilon,v_{i},u_{i}\in[a,b]\}}+\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{v_{i}\notin[a,b]\}}+\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{u_{i}\notin[a,b]\}}.$
As the proof of Lemma B.4, Theorem B.1 indicates that there exists $\delta>0$ and $M>0$, such that for any $m\geq M$,
$\sup_{\begin{subarray}{c}u,v\in[a,b]\\ |u-v|<\delta\end{subarray}}\big|\hat{F}^{-1}_{m}(u)-\hat{F}^{-1}_{m}(v)\big|<\varepsilon$
with probability larger than $1-\varepsilon/4$. Define the event $\mathscr{E}^{{}^{\prime\prime\prime}}$ to be the intersection of this event and the event $\sup_{i\in[n]}|u_{i}-v_{i}|<\delta$. Under the current assumptions of $\hat{Z}_{i}$’s concentration and Lemma B.1, we can see that $\mathbb{P}(\mathscr{E}^{{}^{\prime\prime\prime}})\xrightarrow{m\to\infty}1.$
@@
$u_{i}-G(F(Z_{i})+e_{i})\xrightarrow{\mathbb{P}}0,\quad v_{i}-G(F(Z_{i})+e_{i})\xrightarrow{\mathbb{P}}0.$
Additionally, since $G(F(Z_{i})+e_{i})\sim\text{Uniform}(0,1)$, we have
$\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{v_{i}\notin[a,b]\}}=\mathcal{O}_{\mathbb{P},m,n}(\varepsilon),\quad\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{u_{i}\notin[a,b]\}}=\mathcal{O}_{\mathbb{P},m,n}(\varepsilon).$
Combining these, we have $\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{|\bar{Z}_{i}-T_{i}|\geq\varepsilon\}}=\mathcal{O}_{\mathbb{P},m,n}(\varepsilon)$.
Because $\varepsilon$ is arbitrary, we complete the proof.
@@
$\sup_{(x_{1},x_{2})\in\mathbb{R}^{2}}\big|\bar{F}_{n}(x_{1},x_{2})-F(x_{1},x_{2})\big|\xrightarrow{\mathbb{P}}0\quad\text{as }m,n\to\infty,$
@@
By construction, $T_{i}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}F$ , so by the multivariate Glivenko-Cantelli theorem:
$\sup_{(x_{1},x_{2})\in\mathbb{R}^{2}}\big|F_{n}(x_{1},x_{2})-F(x_{1},x_{2})\big|\xrightarrow{\mathbb{P}}0.$
@@
$\sup_{(x_{1},x_{2})\in\mathbb{R}^{2}}\big|\bar{F}_{n}(x_{1},x_{2})-F_{n}(x_{1},x_{2})\big|\xrightarrow{\mathbb{P}}0.$
For any $(x_{1},x_{2})\in\mathbb{R}^{2}$:
$\big|\bar{F}_{n}(x_{1},x_{2})-F_{n}(x_{1},x_{2})\big|$ $=\left|\frac{1}{n}\sum_{i=1}^{n}\left[\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1},\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1},T_{i2}\leq x_{2}\}}\right]\right|.$